OpenStudy (anonymous):

show all work. find the following product. (x^2-3x+2)(x-4). can someone help me with this problem?

OpenStudy (anonymous):

Can multiply this one? ( x²-3x+2 ) * x = ...

OpenStudy (anonymous):

x^2-(3x)^2+2x

OpenStudy (anonymous):

x^2 * x = .... ? 3x * x = ....?

OpenStudy (anonymous):

Can you try again?

OpenStudy (anonymous):

x3 and 3x^2

OpenStudy (anonymous):

Yes :) Can you put them together now : ( x²-3x+2 ) * x = ...

OpenStudy (anonymous):

x^3-3x^2+2x

OpenStudy (anonymous):

Great =) Now can you multiply with the constant: ( x²-3x+2 ) * ( -4) = .....

OpenStudy (anonymous):

-4x^2+12x-8

OpenStudy (anonymous):

Now again, put YOUR work together =) ( x^3-3x^2+2x ) + ( -4x^2+12x-8 )

OpenStudy (anonymous):

x^3-7x^2+14x-8

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

PERFECT =) See, you did everything by yourself :0

OpenStudy (anonymous):

i just wanted to make sure that i understood this properly

OpenStudy (anonymous):

Well, you're wrong at first; but now you learned and worked it out fine :)

OpenStudy (anonymous):

I mean WAS wrong and have learned :)

OpenStudy (anonymous):

thank god, its just been so long and i still have two other maths to take for my degree

OpenStudy (anonymous):

Well, good luck! As in this problem it's supposedly takes 2 seconds to get it over :)

OpenStudy (anonymous):

thank you for your help